# Behind the Myth of Exploration in Policy Gradients

### v2

@@ 3.2 Illustration of the Effect of Exploration on the Learning Objective
Exploration is of paramount importance in environments with complex dynamics and reward functions, where many locally optimal policies may exist (Lee et al., 2019; Liu & Abbeel, 2021; Zhang et al., 2021b). In the following, we first define such an environment and a policy parameterization that will serve as an example to illustrate the effect of exploration on the optimization process. For the sake of the analysis, we then represent the learning objectives associated to different exploration strategies, and depict their global and local optima. Learning objectives with a single optimum respect the pseudoconcavity criterion. In addition, we represent the neighborhood $\Omega$ of the optimal policy parameters, such that any learning objective with its global maximum within this region is coherent for a given $\epsilon$. In light of the coherence and the pseudoconcavity criteria, we finally elaborate on the policy parameter computed by stochastic gradient ascent algorithms.
We consider the environment illustrated in Figure 1(a) where a car moves in a valley (Bolland et al., 2023). We denote by $x$ and $v$ the position and speed of the car, both composing its state $s=(x,v)$. The valley contains two separate low points, positioned in $x_{initial}=-3$ and $x_{target}=3$, separated by a peak. The car starts at rest $v_{0}=0$ at the highest low point $x_{0}=x_{initial}$ and receives rewards proportional to the depth of the valley at its current position. The reward function is provided in Figure 1(b). We consider a policy $\pi_{K,\sigma}(a|s)=\mathcal{N}(a|\mu_{K}(s),\sigma)$, namely a normally disturbed proportional controller with $\mu_{K}(s)=K\times(x-x_{target})$, parameterized by the vector $\theta=(K,\sigma)$. Figure 1(c) illustrates the contour map of the return of the policy as a function of the parameters $K$ and $\sigma$. The optimal parameters are represented by black dots and correspond to policies that drive the car to pass the peak and reach the lowest valley point in $x_{target}$. The green area represents the set of parameters $\Omega=\left\{\theta^{\prime}|\max_{\theta}J(\pi_{\theta})-J(\pi_{\theta^{\prime}})\leq\epsilon\right\}$ for $\epsilon=1$.
@@
Figure 1: Illustration of the hill environment in Figure 1(a) and its reward function in Figure 1(b). In Figure 1(c), the return of the policy $\pi_{K,\sigma}$ with the global and local maximum represented in black and grey, together with their respective return values.
Figure 2 illustrates the learning objective with the intrinsic rewards $\rho^{s}(s,a)=-\log d^{\pi_{K,\sigma},\gamma}(\phi(s))$, from equation (3), and $\rho^{a}(s,a)=-\log\pi_{K,\sigma}(a|s)$, from equation (4), for different values of the corresponding weights $\lambda_{s}$ and $\lambda_{a}$. Here, the feature is the position in the valley $\phi(s)=x$. First, we observe that for weights approaching zero, the parameter $\theta^{\dagger}$ maximizing the learning objective, represented by a black dot, corresponds to a policy with a high return. More precisely, it is in the green set $\Omega$ such that $\epsilon$-coherence is guaranteed for the value $\epsilon=1$. Larger weights require larger values of $\epsilon$ for guaranteeing the $\epsilon$-coherence criterion. Nevertheless, when increasing the weights, we observe that the learning objective eventually becomes pseudoconcave. There appears to be a trade-off between the two criteria. In Figure 2(b), we observe that in this environment, there is a learning objective that respects the pseudoconcavity criterion and the $\epsilon$-coherence criterion for $\epsilon=1$. Indeed, there is a single global maximum in Figure 2(b) represented by a black dot that is furthermore part of the set $\Omega$.
Shaping the reward function with an exploration strategy based on the state-visitation entropy appears to be a good solution for optimizing the policy. However, a notable drawback is that the reward depends on the policy and its (gradient) computation requires to estimate a complex probability measure. In this example, the intrinsic reward function itself was estimated by Monte-Carlo sampling for every parameter, which would not scale for complex problems and requires approximations and costly evaluation strategies (Islam et al., 2019). In Appendix A we present an alternative problem-dependent intrinsic reward, independent of the policy parameters and thus simple to compute efficiently, that still respects the pseudoconcavity and $\epsilon$-coherence criteria, and in Appendix B we extend the study to more complex environments from the MiniGrid library (Chevalier-Boisvert et al., 2023) where the policy is a deep neural network and the state-visitation probability is approximated.
@@
Figure 2: Contour map of (scaled) learning objective functions for different values of $\lambda_{s}$ and $\lambda_{a}$. The darker the map, the larger the learning objective value. The green area represents the set $\Omega=\left\{\theta^{\prime}|\max_{\theta}J(\pi_{\theta})-J(\pi_{\theta^{\prime}})\leq\epsilon=1\right\}$, such that when the parameter maximizing the learning objective is part of $\Omega$, then the learning objective function is $\epsilon$-coherent with $\epsilon=1$. The black dot is the parameter $\theta^{\dagger}$ globally maximizing the learning objective and the grey dot is the local (non-global) maximum of the learning objective if it exists. Both are labeled with the return values of the corresponding policies.
The observations suggest that well-chosen exploration strategies can lead to learning objective functions that satisfy the two criteria defined in the previous section, thereby guaranteeing that policies suboptimal by at most $\epsilon$ can be computed by local optimization. When designing exploration strategies, it is essential to keep in mind that we modify the learning objective for the algorithms to converge to optimal policy parameters, which can be achieved when both criteria are respected. While strategies such as enforcing entropy can be effective in some environments, they are only heuristic strategies and not to be relied upon exclusively. Furthermore, as illustrated, both criteria may be subject to a trade-off. In more complex environments, an efficient exploration strategy may require to balance both criteria, e.g., through a schedule on the learning objective weights.
@@ 4 Study of the Ascent Direction Distribution
Optimizing pseudoconcave functions with stochastic ascent methods are guaranteed to converge (at a certain rate) under assumptions on the distribution of the gradient estimates at hand (Bottou, 2010; Chen & Luss, 2018; Ajalloeian & Stich, 2020). In this section, we study the influence of the exploration terms on this distribution in the context of policy gradients. More precisely, we study the probability of improving the learning objective and the return with stochastic ascent steps. Intuitively, they shall be sufficiently large for the algorithm to be efficient. We formalize this intuition and illustrate how exploration strategies can increase these probabilities, leading to more efficient algorithms.
@@ 4.1 Policy-Gradient Estimated Ascent Direction
In general, gradient ascent algorithms update parameters in a direction $\hat{d}$ in order to locally improve an objective function $f$. The quality of these algorithms can therefore be studied (for a small step size $\alpha\rightarrow 0$) through the random variable representing the quantity by which the objective increases for each $\theta$
$X$ $=f(\theta+\alpha\hat{d})-f(\theta)=\alpha\>\langle\hat{d},\nabla_{\theta}f(\theta)\rangle\>,$ (7)
[…]
The (asymptotic) convergence of a gradient ascent algorithm is usually studied by bounding the expectation of $X$. Such bounds depend, among others, on the expectation of $\hat{d}$, which equals $\nabla_{\theta}f(\theta)$ when unbiased, and depend on the variance of $\hat{d}$, which deteriorates the expected convergence rate. Finding rates depending on the algorithm is an active field of research. In parallel, a slightly more general problem is to quantify if the expectation of $X$ is driven by rare events. Intuitively, an algorithm with unbiased gradient estimates has positive expected improvements $\mathbb{E}[X]>0$, and should theoretically converge, but may be inefficient in practice if positive events $X>0$ rarely occur. We illustrate in the next section that this phenomenon makes reinforcement learning in sparse-reward environments particularly hard. To the best of our knowledge, no existing result fits to the study of policy gradients. We therefore introduce two new criteria on the probability of improvement $P(X>0)$, which we empirically validate afterwards.
First, we define an exploration strategy as $\delta$-efficient if, and only if, following the ascent direction $\hat{d}\approx\nabla_{\theta}L(\theta)$ has a probability at least $\delta$ to increasing the learning objective $L(\theta)$ almost everywhere. Second, an exploration strategy is $\delta$-attractive if, and only if, there exists a neighborhood of $\theta^{\dagger}$ containing the parameter $\theta^{int}$ maximizing the intrinsic return $J^{int}$, where the probability of increasing the return by following $\hat{d}$ is almost everywhere at least equal to $\delta$. Note that each probability measure and random variable is a function of $\theta$, which we do not explicitly write for the sake of keeping notations simple.
@@
The efficiency criterion quantifies how often a stochastic gradient ascent step improves the learning objective. The larger, the better the learning objective and its stochastic ascent direction approximations. The rationale behind the attraction criterion is that in many exploration strategies, the intrinsic reward is dense, and it is then presumably easy to optimize the intrinsic return in the sense that $\mathbb{P}(\>\langle\hat{i},\nabla_{\theta}J^{int}(\pi_{\theta})\rangle>0)$ is large. It implies that it is easy to locally improve the learning objective by (solely) increasing the value of the intrinsic motivation terms. It furthermore implies that policy-gradient algorithms may be subject to converging towards $\theta^{int}$ rather than $\theta^{\dagger}$ when $\mathbb{P}(\>\langle\hat{d},\nabla_{\theta}J(\pi_{\theta})\rangle>0)$ is small. If the criterion is respected for large $\delta$, the latter is less likely to happen as policy gradients will eventually tend to improve the return of the policy if the parameter approaches $\theta^{int}$ and enters the ball $B(\theta^{\dagger})$; eventually converging towards $\theta^{\dagger}$.
These two new criteria on $\hat{d}$ are independent of the previous ones on $L$, which only captured the quality of the deterministic learning objective functions. In the particular cases where the learning objectives $L$ are $\epsilon$-coherent, for $\epsilon=0$, and pseudoconcave, e.g., with potential-based intrinsic rewards, only the distribution of estimates $\hat{d}$ can explain why some algorithms succeed and others fail. Finally, the value of $\delta$ in the new criteria we introduce can be related to the variance of the estimate $\hat{d}$ under some assumptions, e.g., with Cantelli’s concentration inequalities.
@@ 4.2 Illustration of the Effect of Exploration on the Estimated Ascent Direction
Exploration is usually promoted and tested for problems where the reward function is sparse, typically in maze-environments (Islam et al., 2019; Liu & Abbeel, 2021; Guo et al., 2021). In this section, we first introduce a new maze-environment with sparse rewards where we illustrate the influence of exploration on the gradient estimates of the learning objective. To this end, we present two learning objective functions and elaborate on the influence of exploration on the performance of policy-gradient algorithms in the light of the efficiency and attraction criteria.
Let us consider a maze-environment consisting of a horizontal corridor composed of $S\in\mathbb{N}$ tiles. The state of the environment is the index of the tile $s\in\left\{1,\dots,S\right\}$, and the actions consist in going left $a=-1$ or right $a=+1$. When an action is taken, the agent stays idle with probability $p=0.7$, and moves with probability $0.31-p=$ 0.3 in the direction indicated by the action, then $s^{\prime}=\min(S,\max(1,s+a))$. The agent starts in state $s=1$ and the target state $s=S=15$ is absorbing. Zero rewards are observed except when the agent reaches the target state where a reward $r=100$ is observed. A discount factor of $\gamma=0.99$ is considered. Finally, we study the policy going with probability $\theta$ to the right and probability $1-\theta$ to the left, and with density
$\pi_{\theta}(a|s)=\left\{\begin{array}[]{cl}\theta&\quad\textrm{if }a=1\\ 1-\theta&\quad\textrm{if }a=-1\;.\end{array}\right.$ (13)
The return $J(\pi_{\theta})$ is represented in black in Figure 3(a) as a function of $\theta$ along with two intrinsic returns, $J^{a}(\pi_{\theta})$ in orange and $J^{s}(\pi_{\theta})$ in blue. The intrinsic reward $\rho^{a}(s,a)=-\log\pi_{\theta}(a|s)$, from equation (4), and the intrinsic reward $\rho^{s}(s,a)=-\log d^{\pi_{\theta},\gamma}(s)$, from equation (3), are used respectively. In Figure 3(b), we illustrate the return of the policy without exploration $J(\pi_{\theta})$, along with two learning objective functions, $L^{a}(\theta)$ and $L^{s}(\theta)$, using as exploration strategies the intrinsic returns $J^{a}(\pi_{\theta})$ and $J^{s}(\pi_{\theta})$. We observe that the return is a pseudoconcave function with respect to $\theta$ and the optimal parameter is $\theta^{*}=1$. In addition, the two learning objectives respect the $\epsilon$-coherence criterion for $\epsilon=0$, implying that $\theta^{*}=\theta^{\dagger}$, and respect the pseudoconcavity criterion. It is important to note that with regard to the discussion from Section 3, there is no interest in optimizing the learning objectives rather than directly optimizing the return, as the latter is already pseudoconcave. In the following we illustrate how choosing a correct exploration strategy still deeply influences the policy-gradient algorithms when it comes to building gradient estimates.
@@
Let us compute the estimate $\hat{g}$ and $\hat{d}$ relying on REINFORCE (Williams, 1992) by sampling $8$ histories of length $T=100$. In this particular environment, $\mathbb{P}(D>0)$ equals $\mathbb{P}(G>0)$, and equal the probability that the derivative is positive. We represent in Figure 3(c) this probability for the return and for both learning objectives. First, we see that the learning objectives are more efficient than the return, meaning they are $\delta$-efficient for larger values of $\delta$. Depending on the parameter value, the objective $L^{a}(\theta)$ or $L^{s}(\theta)$ is best in that regard. Second, concerning the attraction criterion, we represent at the top of Figure 3(c) the intervals $B^{a}=[\theta^{int,a},\theta^{\dagger,a}]$ and $B^{s}=[\theta^{int,s},\theta^{\dagger,s}]$. They correspond to the smallest balls containing the maximizers of the intrinsic return and of the learning objective. Let the minima of the orange and blue curves over these intervals be denoted by $\delta^{a}$ and $\delta^{s}$. By definition of the attraction criterion, it is thus respected for any values of $\delta$ at most equal to $\delta^{a}$ and $\delta^{s}$, for $L^{a}(\theta)$ and $L^{s}(\theta)$, respectively. All these observations can eventually be explained as the computation of $\hat{g}$ is always zero when the target is not sampled in the histories, which is highly likely for policies with small values of $\theta$. Policy-gradient algorithms relying on intrinsic exploration would compute optimal policies efficiently where naive optimization without exploration would fail or be sample inefficient.
We have empirically shown that a well-chosen exploration strategy in policy gradients may not only remove local extrema from the objective function, but may also increase the probability that stochastic ascent steps improve the objective function. Under the previous assumptions, this probability measures the efficiency of algorithms. Furthermore, among different learning objectives respecting the coherence and pseudoconcavity criteria, it is best to choose one that has high values for $\delta$ in both the efficiency and attraction criteria. In Appendix A we use these criteria to study other reward-shaping strategies, and in Appendix B we extend the study to more complex environments from the MiniGrid library (Chevalier-Boisvert et al., 2023) where the policy is a deep neural network.
@@
[…]
In the typical case of environments with sparse rewards, many policies observe with high probability state and action pairs with zero rewards and are locally optimal for large probabilities $\Delta$. Typically, in the previous example, the joint set $\{1,\dots,S-2\}\times\{-1,1\}$ is a set of state and action pairs $\mathcal{E}$ that respects the definition equation (14) for large values $\Delta$ when $\theta$ is small. As we have shown, exploration mitigates the convergence of policy-gradient algorithms towards these locally optimal policies. Note that assuming a non-zero reward is uniformly distributed over the state and action space, exploration policies with uniform probabilities over visited states and actions are the best choice for sampling non-zero rewards with high probability. It can thus also be considered as the best choice of exploration to reduce the probability that the stochastic gradient ascent steps do not increase the objective value. Such initial policy may be learned from the framework developed by Lee et al. (2019).

### v3

@@ 3.2 Illustration of the Effect of Exploration on the Learning Objective
Exploration is of paramount importance in environments with complex dynamics and reward functions, where many locally optimal policies may exist (Lee et al.,, 2019; Liu and Abbeel,, 2021; Zhang et al., 2021b, ). In the following, we first define such an environment and a policy parameterization that will serve as an example to illustrate the effect of exploration on the optimization process. For the sake of the analysis, we then represent the learning objectives associated with different exploration strategies, and depict their global and local optima. Learning objectives with a single optimum satisfy the pseudoconcavity criterion. In addition, we represent the neighborhood $\Omega$ of the optimal policy parameters such that any learning objective with its global maximum within this region is coherent for a given $\epsilon$. In light of the coherence and the pseudoconcavity criteria, we finally elaborate on the policy parameters computed by stochastic gradient ascent algorithms.
We consider the environment illustrated in Figure 1(a), where a car moves in a valley (Bolland et al.,, 2023). We denote by $x$ and $v$ the position and speed of the car, which together compose its state $s=(x,v)$. The valley contains two valley floors, located at $x_{initial}=-3$ and $x_{target}=3$, separated by a peak. The car starts at rest $v_{0}=0$ at the higher floor $x_{0}=x_{initial}$ and receives rewards proportional to the depth of the valley at its current position. The reward function is shown in Figure 1(b). We consider a policy $\pi_{K,\sigma}(a|s)=\mathcal{N}(a|\mu_{K}(s),\sigma)$ with $\mu_{K}(s)=K\times(x-x_{target})$, parameterized by the vector $\theta=(K,\sigma)$. It corresponds to a noisy proportional controller. Figure 1(c) illustrates the contour map of the return, annotated with local optima, and shows in green the set of near-optimal policies. The optimal policy drives the car to reach the lowest floor in $x_{target}$.
@@
Figure 1: Illustration of the hill environment in Figure 1(a) and its reward function in Figure 1(b). In Figure 1(c), the return of the policy $\pi_{K,\sigma}$ with the global and local maximum represented in black and grey, together with their respective return values. In green $\Omega=\left\{\theta^{\prime}|\max_{\theta}J(\pi_{\theta})-J(\pi_{\theta^{\prime}})\leq\epsilon=1\right\}$.
Figure 2 illustrates the learning objective with the intrinsic rewards $\rho^{s}(s,a)=-\log d^{\pi_{K,\sigma},\gamma}(\phi(s))$, from equation (4), and $\rho^{a}(s,a)=-\log\pi_{K,\sigma}(a|s)$, from equation (3), for different values of the corresponding weights $\lambda_{s}$ and $\lambda_{a}$. Here, the feature is the position in the valley $\phi(s)=x$. On the one hand, the smaller the exploration weights, the closer the parameter $\theta^{\dagger}$ maximizing the learning objective is to $\theta^{*}$, and the smaller $\epsilon$ for which coherence is guaranteed. Furthermore, the return of the policy corresponding to $\theta^{\dagger}$ is then also higher. In the particular case depicted in the figure, we observe that, only for the smallest weights, the parameter $\theta^{\dagger}$ maximizing the learning objective, represented by a black dot, is inside the green set $\Omega$, such that $\epsilon$-coherence is guaranteed for $\epsilon=1$. On the other hand, as the weights increase, we observe that the learning objective eventually becomes pseudoconcave. As explained, this comes at the cost of a higher $\epsilon$ for guaranteeing $\epsilon$-coherence; there is thus a trade-off between the two criteria. In Figure 2(b), we observe that in this environment, there is a learning objective that satisfies both the pseudoconcavity criterion and the $\epsilon$-coherence criterion for $\epsilon=1$. Indeed, there is a single global maximum represented by a black dot, which is part of $\Omega$.
@@
Figure 2: Contour map of (scaled) learning objective functions for different values of $\lambda_{s}$ and $\lambda_{a}$. The darker the map, the larger the learning objective value. The black dot is the parameter $\theta^{\dagger}$ globally maximizing the learning objective and the grey dot is the local (non-global) maximum of the learning objective if it exists. Both are labeled with the return values of the corresponding policies. The green area represents the set $\Omega=\left\{\theta^{\prime}|\max_{\theta}J(\pi_{\theta})-J(\pi_{\theta^{\prime}})\leq\epsilon=1\right\}$.
Shaping the reward function with an exploration strategy based on the state-visitation entropy appears to be a good solution for optimizing the policy. However, a notable drawback is that the reward depends on the policy, and its (gradient) computation requires estimating a complex probability measure. In this example, the intrinsic reward function itself was estimated by Monte Carlo sampling for each parameter, which would not scale for complex problems and requires approximations and costly evaluation strategies. In Appendix A, we extend the study to more complex environments, where the policy is a neural network and the state-visitation probability is approximated.
The observations suggest that well-chosen exploration strategies can lead to learning objective functions that satisfy the two criteria defined in the previous section, thereby guaranteeing that policies that are suboptimal by at most $\epsilon$ can be computed by local optimization. When designing exploration strategies, it is essential to keep in mind that we modify the learning objective so that the algorithms can converge to optimal policy parameters, which can be achieved when both criteria are satisfied. While strategies such as enforcing entropy can be effective in some environments, they are only heuristic strategies and should not be relied upon exclusively. Furthermore, as illustrated, both criteria may be subject to a trade-off. In more complex environments, an efficient exploration strategy may require balancing both criteria, e.g., through a schedule on the learning objective weights.
@@ 4 Study of the Ascent Direction Distribution
Optimizing pseudoconcave functions with stochastic ascent methods is guaranteed to converge (at a certain rate) under assumptions about the distribution of the gradient estimates (Bottou,, 2010; Chen and Luss,, 2018; Ajalloeian and Stich,, 2020). In this section, we study the influence of the exploration terms on this distribution in the context of policy gradients. More precisely, we study the probabilities of improving the learning objective and the return through stochastic ascent steps. Intuitively, these probabilities should be sufficiently large for the algorithm to be efficient. We formalize this intuition and illustrate how exploration can increase them, leading to more efficient algorithms.
@@ 4.1 Policy-Gradient Estimated Ascent Direction
In general, ascent algorithms update parameters in a direction $\hat{d}$ to locally improve an objective function $f$. The quality of these algorithms can therefore be studied (for a small step size $\alpha\rightarrow 0$) through the random variable representing the amount by which the objective increases for each $\theta$
$X$ $=f(\theta+\alpha\hat{d})-f(\theta)=\alpha\>\langle\hat{d},\nabla_{\theta}f(\theta)\rangle+O(\alpha^{2})\>,$ (7)
where $\langle\cdot,\cdot\rangle$ is the Euclidean scalar product. This variable depends on the random event $\hat{d}$.
The convergence of ascent algorithms is usually studied by bounding the suboptimality of $f(\theta)$ as a function of the number of ascent steps. In this active field of research, the rates are characterized by assumptions on $\hat{d}$ or, equivalently, on $X$. Depending on the assumptions, the efficiency of algorithms is, for example, studied as a function of the statistical moments of the ascent estimate or of the probability of alignment between the gradient and the estimate. We choose to work in the context in which gradient estimates have positive expected improvements $\mathbb{E}[X]>0$ (e.g., because the estimates are unbiased), and we quantify whether this expectation is positive due to rare realizations of large values of $X$. Intuitively, an algorithm with positive expected improvements $\mathbb{E}[X]>0$ should theoretically converge but may be inefficient in practice if positive events $X>0$ rarely occur. Below, we study the influence of such events, typically observed in sparse-reward environments and for which the chosen context is sufficient to assess efficiency without more assumptions. Further work may proceed differently to derive convergence rates or quantify other phenomena.
In the following, we introduce two new criteria on the probability of improvement $P(X>0)$, which we empirically validate later. First, we define an exploration strategy as $\delta$-efficient if, and only if, following the ascent direction $\hat{d}\approx\nabla_{\theta}L(\theta)$ has a probability of at least $\delta$ of increasing the learning objective $L(\theta)$ almost everywhere (within a region of the parameter space). Second, an exploration strategy is $\delta$-attractive if, and only if, there exists a neighborhood of $\theta^{\dagger}$ containing the parameter $\theta^{int}$ that maximizes the intrinsic return $J^{int}$, where the probability of increasing the return by following $\hat{d}$ is almost everywhere at least equal to $\delta$. Note that each probability measure and random variable are functions of $\theta$, which we do not explicitly write to keep the notation simple.
@@
The efficiency criterion quantifies how often a stochastic gradient ascent step improves the learning objective (in a region of the parameter space). The larger, the better the learning objective and its stochastic ascent direction approximations. The rationale behind the attraction criterion is that, in many exploration strategies, the intrinsic reward is dense, and it is then presumably easy to optimize the intrinsic return in the sense that $\mathbb{P}(\>\langle\hat{i},\nabla_{\theta}J^{int}(\pi_{\theta})\rangle>0)$ is large. It implies that it is easy to locally improve the learning objective by (solely) increasing the value of the intrinsic motivation terms. It furthermore implies that policy-gradient algorithms may tend to converge towards $\theta^{int}$ rather than $\theta^{\dagger}$ when $\mathbb{P}(\>\langle\hat{d},\nabla_{\theta}J(\pi_{\theta})\rangle>0)$ is small. If the attraction criterion is satisfied for large $\delta$, the latter is less likely to happen as policy gradients will eventually tend to improve the return of the policy if the parameter approaches $\theta^{int}$ and enters the ball $B(\theta^{\dagger})$; eventually converging towards $\theta^{\dagger}$.
These two new criteria on $\hat{d}$ are independent of the previous ones on $L$, which only captured the quality of the deterministic learning objective functions. In the particular cases where the learning objectives $L$ are $\epsilon$-coherent, for $\epsilon=0$, and pseudoconcave, e.g., with potential-based intrinsic rewards, only the distribution of estimates $\hat{d}$ can explain why some algorithms outperform others. Finally, the value of $\delta$ in the two new criteria could be related to the variance of the estimate $\hat{d}$, for example via Cantelli’s concentration inequalities.
@@ 4.2 Illustration of the Effect of Exploration on the Estimated Ascent Direction
Exploration is usually promoted and tested for problems where the reward function is sparse, typically in maze environments (Islam et al.,, 2019; Liu and Abbeel,, 2021; Guo et al.,, 2021). In this section, we first introduce a new maze environment with sparse rewards to illustrate the influence of exploration on the gradient estimates of the learning objective. To this end, we present two learning objective functions and elaborate on the influence of exploration on the performance of policy-gradient algorithms in light of the efficiency and attraction criteria.
Let us consider a maze environment consisting of a horizontal corridor composed of $S\in\mathbb{N}$ tiles. The state of the environment is the index of the tile $s\in\left\{1,\dots,S\right\}$, and the actions consist of going left a=1a=-1 or right $a=+1$. When an action is taken, the agent stays idle with probability $p=0.7$, and moves with probability 1p=0.31-p=0.3 in the direction indicated by the action, then $s^{\prime}=\min(S,\max(1,s+a))$. The agent starts in state $s=1$, and the target state $s=S=15$ is absorbing. A zero reward is observed except when the agent reaches the target state, where a reward of $r=100$ is observed. A discount factor of $\gamma=0.99$ is considered. Finally, we study the policy that moves right with probability $\theta$ and left with probability $1-\theta$, with density
$\pi_{\theta}(a|s)=\left\{\begin{array}[]{cl}\theta&\quad\textrm{if }a=+1\\ 1-\theta&\quad\textrm{if }a=-1\;.\end{array}\right$. (13)
The return $J(\pi_{\theta})$ is represented in black in Figure 3(a) as a function of $\theta$ along with two intrinsic returns, $J^{a}(\pi_{\theta})$ in orange and $J^{s}(\pi_{\theta})$ in blue. The intrinsic reward $\rho^{a}(s,a)=-\log\pi_{\theta}(a|s)$, from equation (3), and the intrinsic reward $\rho^{s}(s,a)=-\log d^{\pi_{\theta},\gamma}(s)$, from equation (4) are used, respectively. In Figure 3(b), we illustrate the return of the policy without exploration $J(\pi_{\theta})$, along with two learning objective functions, $L^{a}(\theta)$ and $L^{s}(\theta)$, using as exploration strategies the intrinsic returns $J^{a}(\pi_{\theta})$ and $J^{s}(\pi_{\theta})$. We observe that the return is a pseudoconcave function with respect to $\theta$ and the optimal parameter is $\theta^{*}=1$. In addition, the two learning objectives satisfy the $\epsilon$-coherence criterion for $\epsilon=0$, implying that $\theta^{*}=\theta^{\dagger}$, and also satisfy the pseudoconcavity criterion. It is important to note that with regard to the discussion from Section 3, there is no interest in optimizing the learning objectives rather than directly optimizing the return, as the latter is already pseudoconcave. In the following, we illustrate how choosing a correct exploration strategy still deeply influences the policy-gradient algorithms when it comes to building gradient estimates.
@@
Let us compute the estimates $\hat{g}$ and $\hat{d}$ relying on REINFORCE (Williams,, 1992) by sampling $8$ histories of length $T=100$. In this particular environment, $\mathbb{P}(D>0)$ equals $\mathbb{P}(G>0)$, and equals the probability that the derivative is positive. We represent in Figure 3(c) this probability for the return and for both learning objectives. First, we see that the learning objectives are more efficient than the return, meaning they are $\delta$-efficient for larger values of $\delta$. Depending on the parameter value, the objective $L^{a}(\theta)$ or $L^{s}(\theta)$ is best in that regard. Second, concerning the attraction criterion, we represent at the top of Figure 3(c) the intervals $B^{a}=[\theta^{int,a},\theta^{\dagger,a}]$ and $B^{s}=[\theta^{int,s},\theta^{\dagger,s}]$. They correspond to the smallest balls containing the maximizers of the intrinsic return and of the learning objective. Let the minima of the orange and blue curves over these intervals be denoted by $\delta^{a}$ and $\delta^{s}$. By definition of the attraction criterion, it is thus satisfied for any values of $\delta$ at most equal to $\delta^{a}$ and $\delta^{s}$, for $L^{a}(\theta)$ and $L^{s}(\theta)$, respectively. All these observations can eventually be explained because the computation of $\hat{g}$ is always zero when the target is not sampled in the histories, which is highly likely for policies with small values of $\theta$. Policy-gradient algorithms relying on intrinsic exploration would compute optimal policies efficiently where naive optimization without exploration would fail or be sample inefficient.
We have empirically shown that a well-chosen exploration strategy in policy gradients may not only remove local extrema from the objective function, but may also increase the probability that stochastic ascent steps improve the objective function and eventually the return. Under the previous assumptions, this probability measures the efficiency of algorithms. Furthermore, among different learning objectives respecting the coherence and pseudoconcavity criteria, it is best to choose one that has high values for $\delta$ in both the efficiency and attraction criteria. In Appendix A, we extend the study to more complex environments and policies.
@@
$\exists\mathcal{E}$ $\in\left\{\mathcal{X}\Big{|}\underset{\begin{subarray}{c}s\sim d^{\pi,\gamma}(\cdot)\\ a\sim\pi(\cdot|s)\end{subarray}}{\mathbb{E}}\big{[}\mathbbm{1}\left[(s,a)\in\mathcal{X}\right]\big{]}\geq\Delta\right\}:\pi\in\arg\max_{\pi^{\prime}}\underset{\begin{subarray}{c}s\sim d^{\pi^{\prime},\gamma}(\cdot)\\ a\sim\pi^{\prime}(\cdot|s)\end{subarray}}{\mathbb{E}}\big{[}\mathbbm{1}\left[(s,a)\in\mathcal{E}\right]\rho(s,a)\big{]}\;.$ (14)
In the case of sparse rewards, many policies observe with high probability state and action pairs with zero rewards and are locally optimal for large probabilities $\Delta$. Typically, in the previous example, the joint set $\{1,\dots,S-2\}\times\{-1,1\}$ is a set of state and action pairs $\mathcal{E}$ that satisfies the definition in equation (14) for large values $\Delta$ when $\theta$ is small. As we have shown, exploration mitigates the convergence of policy-gradient algorithms towards these locally optimal policies. Note that assuming a non-zero reward is uniformly distributed over the state and action space, exploration policies with uniform probabilities over visited states and actions are the best choice for sampling non-zero rewards with high probability. It can thus also be considered the best choice of exploration to reduce the probability that the stochastic gradient ascent steps do not increase the objective value. Such initial policy may be learned as suggested by Lee et al., (2019).